OpenStudy (anonymous):

A certain starship can fly 952 miles with the wind in 4 hours and travel the same distance against the wind in 7 hours. What is the speed of the wind?

OpenStudy (anonymous):

r t = d Solve the following for r and w, speed and wind respectively.\[\{(r+w)4=952,(r-w)7\text{=}952\}\]\[\{r=187 \text{ mph},w=51 \text{ mph}\}\]